OpenStudy (anonymous):

If BA=13, BC=2x-3 and AD=5, then find DC.

OpenStudy (anonymous):

BA+BC=AC=10+2x AD+DC=AC 5+DC=10+2x DC=5+2x

OpenStudy (anonymous):

How do I know that Ba+bc=ac... is this because of how the triangle is drawn?

OpenStudy (anonymous):

 do you understand what i mean?